# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ 3. Probabilistic Setup for Aggregated and Sampled Dynamics
This section introduces a rigorous probabilistic framework for the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
@@ 3.1. Aggregated dynamics
Fix a probability space $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ on which a $d$-dimensional Brownian motion $\{W_{t}\}_{t\geq 0}$ is defined, and let ${\mathbb{F}}^{W}\coloneqq\{\mathcal{F}_{t}^{W}\}_{t\geq 0}$ be the $\mathbb{P}^{W}$-completion of the filtration generated by $W$, which satisfies the usual conditions of completeness and right-continuity. Throughout this paper, we impose the following regularity condition on the action space $A$, the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
@@
    $({A},d_{A})$ is a Polish space equipped with the Borel $\sigma$-algebra $\mathcal{B}({A})$, and $a_{0}\in A$;
  2. (2)
    $b:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d\times d}$ are measurable functions. There exists $C\geq 0$ such that for all $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
@@
    $\bm{\pi}:[0,T]\times\mathbb{R}^{d}\times\mathcal{B}({A})\to[0,1]$ is a probability kernel. For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$.
@@
Here, we focus on strong solutions of (2.4) rather than weak solutions, which allows us to identify $W$ as the only driving noise in the state process $\tilde{X}$. This simplifies the construction of the sampled dynamics and the subsequent error estimates in terms of the sampling frequency. Under (H.1), Condition (H.2) holds if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are locally Lipschitz continuous in the state variable (cf. [27, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [19, Lemma 2]). Some typical sufficient conditions on $b,\sigma$ and $\bm{\pi}$ to guarantee these regularity conditions on $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are given in Examples 4.1, 4.2 and 4.3.
@@ 3.2. Sampled dynamics with randomized actions
As $A$ is a Polish space, a sampling procedure of a probability kernel $\bm{\pi}$ always exists. Indeed, by [21, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$. This approach has been utilized in [11, 4, 3]. However, the corresponding function $\phi$ is guaranteed to exist only through a measure-theoretic argument, making it difficult to implement such a sampling procedure.
@@
which will be referred to as the sampled dynamics [^2] This definition slightly differs from the definition of the “grid sampling SDE” in [3, Section 3] even though both rely on drawing finitely many samples on the grid point. In our definition, the action sequence $a_{t}$ is a piecewise constant and is given by a sampling function at the grid points $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$. By contrast, in the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant. in the subsequent analysis. For notational convenience, we write (3.2) in the following equivalent form:
@@
Note that the sampled dynamics evolve continuously over time while the control process remains constant within each subinterval. In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $[t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$. This aligns with the RL setting, where the agent observes a continuous-time system at discrete time points and adjusts their actions based on these observations.
@@
Suppose (H.1) holds. Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ be the $\mathbb{P}$-null set. Then for any grid $\mathscr{G}$, (3.3) admits a unique ${\mathbb{F}}$-adapted strong solution $X^{\mathscr{G}}$ on the space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, for all $p\geq 2$, there exists a constant $C_{p}\geq 0$ such that for all grids $\mathscr{G}$ and all $t\in[0,T]$, $\mathbb{E}\left[|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}t}$, and $X^{\mathscr{G}}_{t}$ is adapted to $\mathcal{G}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{t_{i}<t}\}\vee\mathcal{N}$.
@@ 4.1.1. Weak convergence for regular coefficients
To highlight the key ideas without introducing needless technicalities, we begin by providing the analysis for sufficiently smooth coefficients and test functions. In this setting, we obtain the maximum convergence rate, which is of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero within the regular time grids $\mathscr{P}_{[0,T]}$.
To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p>0$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
@@
    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$;
@@
Condition (H.3) holds if the coefficients of the aggregated dynamics (2.4) are sufficiently regular. For instance, if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are in the class $C_{p}^{4}$ and have bounded derivatives, then (H.3) can be verified using the probabilistic argument in [28].
@@
$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
@@
For a given time grid $\mathscr{G}$, we execute the policy only at the nodes $t_{i}$, and use the same action over the time interval $[t_{i},t_{i+1})$ to obtain the sampled dynamics $X^{\mathscr{G}}$. Theorem 4.1 illustrates that with this policy execution method, for a given test function $f$, the error between $\mathbb{E}[f(X^{\mathscr{G}}_{t})]$ and the original aggregated dynamics $\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ converges to 00 at the optimal first-order convergence $O(|\mathscr{G}|)$.
@@ 4.1.2. Weak convergence with irregular coefficients
For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index $l-[l]$ in $x$ and with index $(l-2r-|s|)/2$ in $t$, i.e.,
$\|u\|_{T}^{(l)}:=$ $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
@@
Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [24, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
@@ 4.2. Weak error conditioned on execution noise
Compared with (H.3), (H.5) imposes the additional condition on the uniform boundedness $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, which allows for controlling the concentration behavior of the sampled dynamics with respect to the execution noise $\xi$. It is easy to see that (H.4) implies (H.5).
The following theorem shows that the weak convergence rate in expectation, established in Section 4.1, also holds in a high probability sense with respect to the sampling noises. It extends a similar result established in [32] for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
@@
Suppose (H.1), (H.2), (H.3) and (H.5) hold. Then for the test function $f$ in Theorem 4.1, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho.$ (4.15)
Assume further that (H.1), (H.2) and (H.4) hold. Then for the test function $f$ in Theorem 4.2, there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho,$ (4.16)

### v2

@@ 3. Probabilistic Setup for Aggregated and Sampled Dynamics
This section rigorously formulates the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
@@ 3.1. Aggregated dynamics
Fix a probability space $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ on which a $d$-dimensional Brownian motion $\{W_{t}\}_{t\geq 0}$ is defined, and let ${\mathbb{F}}^{W}\coloneqq\{\mathcal{F}_{t}^{W}\}_{t\geq 0}$ be the $\mathbb{P}^{W}$-completion of the filtration generated by $W$, which satisfies the usual conditions of completeness and right-continuity. Let the action set $A$ be a nonempty Polish space equipped with the metric $d_{A}$, and let $\mathcal{B}({A})$ be the Borel $\sigma$-algebra on $A$. Throughout this paper, we impose the following regularity condition on the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
@@
    $b:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d\times d}$ are measurable functions. There exists $C\geq 0$ such that for a fixed $a_{0}\in A$, any $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
@@
    $\bm{\pi}:[0,T]\times\mathbb{R}^{d}\times\mathcal{B}({A})\to[0,1]$ is a probability kernel. For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$, with a fixed $a_{0}\in A$.
@@
Here, we focus on strong solutions of (2.4) rather than weak solutions, which allows us to identify $W$ as the only driving noise in the state process $\tilde{X}$. This simplifies the construction of the sampled dynamics and the subsequent error estimates in terms of the sampling frequency. Under (H.1), Condition (H.2) holds if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are locally Lipschitz continuous in the state variable (cf. [31, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [21, Lemma 2]). Some typical sufficient conditions on $b,\sigma$ and $\bm{\pi}$ to guarantee these regularity conditions on $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are given in Examples 4.1, 4.2 and 4.3.
@@ 3.2. Sampled dynamics with randomized actions
As $A$ is a Polish space, a sampling procedure of a probability kernel $\bm{\pi}$ always exists. Indeed, by [23, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$. This approach has been utilized in [12, 4, 3]. However, the corresponding function $\phi$ is guaranteed to exist only through a measure-theoretic argument, making it difficult to implement such a sampling procedure.
@@
which will be referred to as the sampled dynamics [^4] This definition slightly differs from the definition of the “grid sampling SDE” in [3, Section 3] even though both rely on drawing finitely many samples on the grid point. In our definition, the action sequence $a_{t}$ is a piecewise constant and is given by a sampling function at the grid points $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$. In contrast, for the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant. in the subsequent analysis. For notational convenience, we write (3.2) in the following equivalent form:
@@
Note that the sampled dynamics evolve continuously over time while the control process remains constant within each subinterval. In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$. This aligns with the RL setting, where the agent observes a continuous-time system at discrete time points and adjusts their actions based on these observations.
@@
Suppose (H.1) holds. Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ is the $\sigma$-algebra of $\mathbb{P}$-null sets. Then for any grid $\mathscr{G}$, (3.3) admits a unique $\mathbb{F}$}-adapted strong solution $X^{\mathscr{G}}$ on the space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, for all $p\geq 2$, there exists a constant $C_{p}\geq 0$ such that for all grids $\mathscr{G}$ and all $t\in[0,T]$, $\mathbb{E}\left[|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}t}$, and $X^{\mathscr{G}}_{t}$ is adapted to $\mathcal{G}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{t_{i}<t}\}\vee\mathcal{N}$.
@@ 4.1.1. Weak convergence for regular coefficients
To highlight the key ideas without introducing needless technicalities, we begin by providing the analysis for sufficiently smooth coefficients and test functions. In this setting, we obtain the maximum convergence rate of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero.
To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p\in{\mathbb{N}}_{0}$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
@@
    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$ for some $p\in{\mathbb{N}}_{0}$;
@@
To obtain maximal first-order weak convergence, Condition (H.3) requires that both the state coefficients and the solution to the associated Kolmogorov PDE possess sufficient regularity. This regularity requirement is stronger than Condition (H.6) assumed for half-order strong convergence. This distinction is consistent with the fact that establishing first-order weak convergence of Euler–Maruyama schemes for SDEs typically requires stronger regularity conditions than the Lipschitz continuity that ensures strong convergence (see, e.g., [25, Theorem 14.1.5] and [17] for the conditions of weak and strong convergence of Euler–type schemes, respectively). However, as emphasized in Section 1, our convergence result does not follow from existing results on Euler–Maruyama schemes, since the sampled dynamics (3.3) have random coefficients and evolve continuously, and consequently do not arise from a standard Euler–Maruyama discretization of the aggregated dynamics (2.4).
The PDE solution regularity in Condition (H.3) holds when the coefficients of the aggregated dynamics (2.4) are sufficiently regular. For instance, if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are in the class $C_{p}^{4}$ and have bounded derivatives, then (H.3) can be verified using the probabilistic argument in [33]. Alternatively, if the diffusion matrix $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic and $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ are bounded functions in $C^{3}_{p}$ with bounded derivatives, then the desired solution regularity follows from the standard regularity theory of parabolic PDEs; see e.g., [30, Lemma 2.2] and [27, Chapter 4, Theorem 5.1]. The details are left to the interested reader.
The following theorem establishes the optimal first-order weak convergence of the sampled dynamics to the aggregated dynamics.
@@
$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
@@
By Theorem 4.1, for a given function $f\in C_{p}^{4}(\mathbb{R}^{d})$, $f(X^{\mathscr{G}}_{T})$ is a biased estimator of $\mathbb{E}[f(\widetilde{X}_{T})]$, and the bias vanishes at first order, as the sampling mesh size $|\mathscr{G}|$ tends to zero. Moreover, by Lemma 3.1, the variance of the estimator $f(X^{\mathscr{G}}_{T})$ is uniformly bounded, independent of the mesh size $|\mathscr{G}|$. The computational cost of generating one trajectory of $X^{\mathscr{G}}_{T}$ increases linearly with the number of grid points in $\mathscr{G}$. For further details on its implications for the sample complexity of Monte Carlo estimators of $\mathbb{E}[f(\widetilde{X}_{t})]$, we refer to Remark 4.3.
@@ 4.1.2. Weak convergence with irregular coefficients
For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to{\color[rgb]{0,0,0}E}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index l[l]l-[l] in $x$ and with index (l2r|s|)/2(l-2r-|s|)/2 in $t$, i.e.,
$\|u\|_{T}^{(l)}$= $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
@@
Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [27, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
@@ 4.2. Weak error conditioned on execution noise
Compared with (H.3), (H.5) imposes the additional condition on the uniform boundedness $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, which allows for controlling the concentration behavior of the sampled dynamics with respect to the execution noise $\xi$. It is easy to see that (H.4) includes the uniform boundedness assumption on $a$, thus, if (H.4) holds, then we do not need (H.5).
The following theorem establishes that the sampled dynamics converge weakly at half-order, averaged over the system noise, uniformly over the realizations of the sampling noises with high probability. It extends [37, Theorem 2.2] established for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
@@
Suppose (H.1), (H.2), (H.3) and (H.5) hold. Then for all $f\in C^{4}_{p}({\mathbb{R}}^{d})$, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big(1+\sqrt{\ln(2/\varrho)}\big)\right]\leq\varrho.$ (4.15)
Suppose instead that (H.1), (H.2) and (H.4) hold. Then for all $f\in\mathcal{H}^{(l+2)}$ with $l$ in (H.4), there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big(1+\sqrt{\ln(2/\varrho)}\big)\right]\leq\varrho,$ (4.16)
@@
Note that Theorem 4.3 shows that the convergence of the sampled dynamics to the aggregated dynamics can be achieved simply by increasing the sampling frequency, without the need to average over multiple realizations of the sampling noise at fixed time points. This observation significantly reduces the sample complexity of estimators for the aggregated dynamics.
To see it, consider estimating $\mathbb{E}[f(\widetilde{X}_{T})]$ using the sampled dynamics (3.2). Let $\mathscr{G}_{n}=\{i\frac{T}{n}\}_{i=0}^{n}$, $n\in{\mathbb{N}}$, be a uniform time grid, and let $(W_{k})_{k\in{\mathbb{N}}}$ be a collection of independent $d$-dimensional Brownian motions, which are also independent of $\xi$. For each $m\in{\mathbb{N}}$, define the estimator $\hat{I}_{m,n}$ of $\mathbb{E}[f(\widetilde{X}_{T})]$ by
$\hat{I}_{m,n}\coloneqq\frac{1}{m}\sum_{k=1}^{m}f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)),$ (4.17)
where $X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)$ denotes the terminal state of the sampled dynamics (3.2), driven by the Brownian motion $W_{k}$ and the common sampling noises $\xi$ on the grid $\mathscr{G}_{n}$. We define the complexity of $\hat{I}_{m,n}$ as the total number of random variables, including the Brownian motions $(W_{k})_{k\in{\mathbb{N}}}$ and the sampling noises $(\xi_{n})_{n\in{\mathbb{N}}_{0}}$, required to evaluate a single realization of $\hat{I}_{m,n}$.
Corollary 4.1.
Suppose (H.1), (H.2), (H.3) and (H.5) hold. Let $f\in C^{4}_{p}({\mathbb{R}}^{d})$ be bounded. There is a constant $C\geq 0$ such that for all $\varepsilon\in(0,1)$, by setting $m_{\varepsilon}=n_{\varepsilon}=\lceil\varepsilon^{-2}\left(1+\ln(2/\varepsilon)\right)\rceil$, it holds with probability at least $1-\varepsilon$ that $|\hat{I}_{m_{\varepsilon},n_{\varepsilon}}-\mathbb{E}[f(\widetilde{X}_{T})]|\leq C\varepsilon$. Moreover, the sample complexity of $\hat{I}_{m_{\varepsilon},n_{\varepsilon}}$ is $2\lceil\varepsilon^{-2}\left(1+\ln(2/\varepsilon)\right)\rceil$.
The proof of Corollary 4.1 is given in Section 6.2. Similar results hold when the objective involves a sufficiently regular running reward of the aggregated process, as in the cases analyzed in Section 5.
Remark 4.3.
Note that in the idealized scenario where one can directly sample trajectories of the aggregated dynamics, a Monte Carlo estimator of $\mathbb{E}[f(\widetilde{X}_{T})]$ achieves a sample complexity of order $\mathcal{O}(\varepsilon^{-2})$. As shown in Corollary 4.1, the estimator (4.17) attains this optimal sample complexity up to a logarithmic factor. This is due to the fact that the estimator (4.17) does not regenerate sampling noises for different Brownian motion trajectories.
In contrast, the weak convergence results in Theorems 4.1 and 4.2 suggest considering the naive Monte Carlo estimator
$\tilde{I}_{m,n}\coloneqq\frac{1}{m}\sum_{k=1}^{m}f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi_{k})),$ (4.18)
where $(\xi_{k})_{k\in{\mathbb{N}}}$ are independent copies of the sampling noises $\xi$. Under the same conditions as in Corollary 4.1, the standard bias–variance decomposition shows that for all $m,n\in{\mathbb{N}}$,
${\mathbb{E}}[|\tilde{I}_{m,n}-\mathbb{E}[f(\widetilde{X}_{T})]|^{2}]^{1/2}\leq C\left(\frac{1}{\sqrt{m}}+\frac{1}{n}\right),$
for a constant $C\geq 0$ independent of $m$ and $n$. To achieve a total error $\varepsilon$, we choose $m_{\varepsilon}=\mathcal{O}(\varepsilon^{-2})$ and $n_{\varepsilon}=\mathcal{O}(\varepsilon^{-1})$. However, since the estimator (4.18) requires regenerating sampling noises for each trajectory, the overall complexity is $\mathcal{O}(m_{\varepsilon}n_{\varepsilon})=\mathcal{O}(\varepsilon^{-3})$, which is significantly higher than that of the estimator (4.17).